### v2

# Higher-order Gini indices: An axiomatic approach

Abstract
Via an axiomatic approach, we characterize the family of $n$-th order Gini deviation, defined as the expected range over $n$ independent draws from a distribution, to quantify joint dispersion across multiple observations. This family extends the classical Gini deviation, which relies solely on pairwise comparisons. The normalized version is called a high-order Gini coefficient. The generalized indices grow increasingly sensitive to tail inequality as $n$ increases, offering a more nuanced view of distributional extremes. The higher-order Gini deviations admit a Choquet integral representation, inheriting the desirable properties of coherent deviation measures. Furthermore, we show that both the $n$-th order Gini deviation and the $n$-th order Gini coefficient are statistically $n$-observation elicitable, allowing for direct computation through empirical risk minimization. Data analysis using World Inequality Database data reveals that higher-order Gini coefficients capture disparities that the classical Gini coefficient may fail to reflect, particularly in cases of extreme income or wealth concentration.
Keywords: Gini coefficient, elicitability, inequality measurement, Choquet integrals, risk measures
1 Introduction
The Gini deviation (GD) and Gini coefficient (abbreviated as GC, also called the Gini index) are fundamental tools for measuring inequality and dispersion in economic research, owing to its intuitive interpretation, broad comparability, and robust statistical properties. Originally introduced by Gini (1912, 1921), the classical GC is defined through its geometric definition based on the Lorenz curve. Its intuitive interpretation and straightforward quantile representation have established it as a standard summary statistic (Gastwirth (1972) and Lerman and Yitzhaki (1984)). Beyond its theoretical appeal, GC is widely used in practice to summarize and compare income or wealth distributions across countries and over time (Kopczuk et al. (2010)). Policymakers rely on it to monitor inequality trends, evaluate the effects of tax and transfer systems, and design social welfare programs (Duclos (2000) and Kenworthy and Pontusson (2005)). For example, an increase in GC is often interpreted as a signal of rising income concentration, motivating debates over progressive taxation or redistribution. In international development, organizations such as the World Bank and the Organisation for Economic Co-operation and Development use the Gini index as a headline indicator to track progress toward reducing inequality and promoting inclusive growth. The measurement of economic inequality has been extensively treated in numerous monographs and handbooks, with GC and its various extensions occupying a central position in this literature; see, e.g., Atkinson and Piketty (2007), Banerjee and Duflo (2011) and Atkinson and Bourguignon (2014). Beyond public policy, Gini-type indices, especially GD, are employed in actuarial risk measurement (Denneberg (1990), Furman and Zitikis (2017) and Furman et al. (2017)), decision analysis (Eeckhoudt and Laeven (2021), who called GD as maxiance), portfolio optimization (Yitzhaki (1982) and Ringuest et al. (2004)), and statistical scoring rules (Zheng et al. (2000) and Gneiting (2011)).
Both GD and GC can be formulated as functions of random variables. For a random variable $X$ that represents a distribution of wealth or income, GD averages relative dispersion through pairwise differences:
$\mathrm{GD}(X)=\frac{1}{2}\,\mathbb{E}[|X-X^{\prime}|],$ (1)
where $X^{\prime}$ is an independent and identically distributed (iid) copy of $X$. The GC of $X$ is defined as
$\mathrm{GC}(X)=\frac{\mathrm{GD}(X)}{\mathbb{E}[X]}=\frac{\mathbb{E}[|X-X^{\prime}|]}{\mathbb{E}[X+X^{\prime}]},$
where we assume $X\geq 0$ and $\mathbb{E}[X]>0$. Both GD and GC use two iid observations to capture the inequality via all possible pairwise contrasts in the distribution.
Figure 1: Values of $\mathrm{GC}$ based on household wealth for the United States, China, Canada, and the United Kingdom from 2000 to 2023
Despite its popularity, GC is only one measure of distributional inequality or dispersion, and it may fail to report some practically important features of the underlying distribution. [^1] The measurement of inequality is a complex task. For instance, Clarke and Kopczuk (2025) provided a comprehensive discussion on various concepts and definitions in measuring income inequality. Our paper focuses on the choice of the indices as measurement tool, implicitly assuming that the quantity to be measured (e.g., income or wealth) is well defined. Figure 1 presents the GC curves of wealth distributions for four economies—US, China (mainland), UK, and Canada—over the period of 2000 to 2023, with data from the World Inequality Database (WID). [^2] WID (http://wid.world) provides the full wealth distribution for each country, but not the full income distribution for each country. Therefore, we take the wealth distributions as the main example. As we can see from the figure, US has the largest GC among the four countries, and UK, Canada and China have similar GC curves since 2010, although the GC curve of China is substantially smaller than the others before 2010. From these curves, one may be tempted to make the following conclusions.
  1. 1.
    China has a similar wealth distribution to Canada and UK since 2010.
  2. 2.
    The wealth distribution in China has a similar level of equality to Canada and UK since 2010.
  3. 3.
    The wealth inequality of China has been increasing up to 2010, and it is stabilized after 2010.
To better understand these statements, we plot in Figure 2 the wealth shares held by the top 1% and top 10% of the population in China, Canada, and UK. The data clearly show that both the top 1% and top 10% populations in China have consistently held a larger share of total wealth than their counterparts in Canada and the UK after 2010. This provides evidence against the first two statements, indicating that China’s wealth distribution has been notably more unequal than UK and Canada. On the other hand, both Figures 1 and 2 agree that the wealth distribution in China may be stabilized after 2010.
Figure 2: Wealth shares held by the top 1% (left) and top 10% (right) in the United States, China, Canada, and the United Kingdom from 2000 to 2023.
The above discussion suggests that GC may not provide the most relevant information about wealth inequality, although it is the most popular index for this purpose. To address this limitation of GC, we consider generalizing GC to a family of indices. We begin by examining key properties satisfied by GD and GC, which we hope to keep. A key insight of GD in (1) is that it can be expressed as the expected value of a function of two independent copies of the underlying variable. Motivated by this observation, we consider natural extensions that involve expectations of functions of $n$ iid copies, with $n\geq 2$. Building on this idea, together with several other natural properties satisfied by GD—including symmetry, comonotonic additivity, and continuity—we develop an axiomatic characterization showing that any measure satisfying these axioms can be expressed as a linear combination of functionals in the following family, which we call the higher-order Gini deviations,
$\mathrm{GD}_{n}(X)=\frac{1}{n}\,\mathbb{E}\bigl{[}\max\{X_{1},\dots,X_{n}\}-\min\{X_{1},\dots,X_{n}\}\bigr{]},$
where $X_{1},\dots,X_{n}$ are iid copies of $X$. The proof of our axiomatic characterization goes by first establishing a signed Choquet integral representation and then identifying those that can be represented by expected values of functions of iid observations. Many classes of decision models, such as the dual utility of Yaari (1987) and the Choquet expected utility model of Schmeidler (1989), are formulated based on Choquet integrals; see the more recent works of Strack and Wambach (2017), Gershkov et al. (2025) and Hartmann and Kauffeldt (2025). The main differences here is that our functionals $\mathrm{GD}_{n}$ do not satisfy monotonicity, thus they are called “signed”.
Observing from (1) that the Gini deviation admits the alternative representation
$\mathrm{GD}(X)=\frac{1}{2}\mathbb{E}[\max\{X,X^{\prime}\}-\min\{X,X^{\prime}\}],$
it follows that $\mathrm{GD}_{n}$ with $n=2$ coincides with GD. Thus, the $n$-th order Gini deviation $\mathrm{GD}_{n}$ replaces the pairwise absolute difference with the expected range of $n$ independent draws, thus quantifying the average spread within larger groups. As $n$ increases, $\mathrm{GD}_{n}$ becomes increasingly sensitive to tail behavior and concentration, since the range over $n$ iid draws amplifies the impact of extreme values. Based on $\mathrm{GD}_{n}$, the corresponding $n$-th order Gini coefficient is defined by
$\mathrm{GC}_{n}(X)=\frac{\mathrm{GD}_{n}(X)}{\mathbb{E}[X]}.$
The quantity $\mathrm{GC}_{n}$ as a generalization of GC was recently proposed by Gavilan-Ruiz et al. (2024), but a comprehensive study is still lacking, particularly with respect to an axiomatic framework and empirical analysis.
Figure 3 depicts $\mathrm{GC}_{10}$ for the four countries. The values of $\mathrm{GC}_{10}$ in China increase rapidly during the considered period, surpassing those of Canada and UK, and gradually approaching the levels observed in US. This upward trend underscores the ability of $\mathrm{GC}_{10}$ to capture shifts in upper-tail wealth concentration, and it is consistent with the observations from Figure 2 regarding statements 1–3 above. Therefore, $\mathrm{GC}_{n}$ provides interpretable information that is complementary to the classic $\mathrm{GC}$, better capturing more nuanced features of the wealth distributions. A more comprehensive analysis in Section 7 illustrates $\mathrm{GC}_{n}$ with other values of $n$ and other datasets.
Figure 3: Values of $\mathrm{GC}_{10}$ based on household wealth for the United States, China, Canada, and the United Kingdom from 2000 to 2023
Many other variants of GD and GC exist in the literature; see Ceriani and Verme (2012) for several formulations of GC proposed by Gini, and Yitzhaki (1998, 2003), Yitzhaki and Schechtman (2012) and Vila and Saulo (2025) for further developments.
After reviewing the classical Gini deviation and Gini coefficient in Section 2, we first present the axioms satisfied by the Gini indices and then develop a rigorous axiomatic framework for their higher-order generalizations in Section 3. In Section 4, we explore the key properties of the higher-order Gini indices in detail. First, we examine the economic properties of the higher-order Gini coefficient, emphasizing its interpretability and practical relevance. Second, we show that $\mathrm{GD}_{n}$ admits a representation as a signed Choquet integral with a concave distortion function. Moreover, $\mathrm{GD}_{n}$ belongs to the class of deviation measures in the sense of Rockafellar et al. (2006) and the class of coherent measures of variability in the sense of Furman et al. (2017). This integral formulation guarantees several desirable properties, including symmetry, convexity, and comonotonic additivity, and leads to economically useful properties of $\mathrm{GC}_{n}$, such as quasi-convexity and mixture-quasi-concavity. Finally, we present a comparative analysis between the $n$-th order GD and the standard deviation.
We study the multi-observation elicitability of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ in Section 5. Elicitability refers to the property of a statistical functional being the unique minimizer of an expected scoring function, which is central to forecast evaluation and model comparison (Gneiting (2011) and Frongillo and Kash (2021)). Many standard risk measures, including general deviation measures, are not elicitable in the classical (single-observation) framework (Wang and Wei (2020)). To address this, the notion of multi-observation elicitability—where the score function depends on multiple independent observations—was introduced by Casalaina-Martin et al. (2017); see also Frongillo et al. (2017). Following from their definitions as the expected value of functions of iid observations or their ratio, we show that both $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ are $n$-observation elicitable, thereby enabling rigorous comparative backtesting in risk and inequality measurement contexts. Section 6 contains results on the statistical inference of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ using its standard empirical estimators, including their consistency and asymptotic variance.
We do not intend to argue that the higher-order Gini indices $\mathrm{GC}_{n}$ necessarily outperform the classical case $\mathrm{GC}$ in all settings. Instead, we argue that higher-order indices can uncover features that $\mathrm{GC}$ fails to detect, thus providing complementary information, as explained in the example above. We further provide more comprehensive empirical analyses using data from WID in Section 7, spanning post-tax national income and household wealth across countries and regions between 2000 and 2023. We find that higher-order Gini indices such as $\mathrm{GC}_{5}$ or $\mathrm{GC}_{10}$ provide additional insights in settings where $\mathrm{GC}$ fails to distinguish some distributions. These findings underscore the value of $\mathrm{GC}_{n}$ as a complementary tool for inequality assessment, especially in the context of evaluating tax policies, monitoring ultra-high-net-worth individuals, or studying long-term distributional trends.
Finally, Section 8 concludes the paper. All proofs are put into Appendix A. Appendix B contains explicit expressions for the higher-order Gini indices under several commonly used parametric distributions. The monotonicity properties of ${\rm GD}_{n}$ and ${\rm GC}_{n}$ with respect to the order $n$, along with simulation results of comparisons across different distributions and orders, are presented in Appendix C.
2 The Gini deviation and Gini coefficient
We fix an atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$. For $q\in[1,\infty)$, let $L^{q}$ be the set of random variables with a finite $q$th moment, and $L^{\infty}$ is the set of essentially bounded random variables. Almost surely equal random variables are treated as identical. For a random variable $X$, we use $F_{X}$ for its distribution function, and let $F^{-1}_{X}(t)$ be its left $t$-quantile, given by $F^{-1}_{X}(t)=\inf\{x\in\mathbb{R}:\mathbb{P}(X\leq x)\geq t\}$ for $t\in(0,1)$. Let $\Delta_{n}$ denote the standard $n$-simplex, that is, $\Delta_{n}=\left\{(x_{1},\dots,x_{n})\in[0,1]^{n}:x_{1}+\dots+x_{n}=1\right\}.$
Let $\mathcal{M}^{q}$ be the distributions of elements of $L^{q}$ for $q\in[1,\infty]$. Let $L^{q}_{+}=\{X\in L^{q}:X\geq 0,~X\neq 0\}$ and $\mathcal{M}^{q}_{+}$ be the distributions of elements of $L^{q}_{+}$. For a set $\mathcal{X}$ of random variables and a mapping $\rho:\mathcal{X}\to\mathbb{R}^{k}$ that is law invariant (i.e., $\rho(X)=\rho(Y)$ if $X$ and $Y$ are identically distributed), we abuse the notation by treating it also as a mapping from a set $\mathcal{M}_{\rho}$ of distributions to $\mathbb{R}^{k}$, that is, $\rho(F_{X})=\rho(X)$. In this case, $\mathcal{M}_{\rho}$ is the set of distributions of random variables in $\mathcal{X}$. Since we focus on law-invariant mappings throughout the paper, this abuse of notation will be helpful in many places.
The Gini deviation (GD) and the Gini coefficient (GC) are two very important indices of dispersion, deviation, and economic inequality, and they have wide applications in finance, economics and machine learning. The Gini deviation $\mathrm{GD}:L^{1}\to\mathbb{R}$ is defined as
$\mathrm{GD}(X)=\frac{1}{2}\mathbb{E}[|X-X^{\prime}|],$ (2)
where $X^{\prime}$ is an iid copy of $X$. Alternatively, we can represent $\mathrm{GD}$ in terms of quantile functions or distribution functions, via (see e.g., Wang et al. (2020b, Example 1))
$\mathrm{GD}(X)=\int_{0}^{1}F_{X}^{-1}(t)(2t-1)\,\mathrm{d}t=\int_{\mathbb{R}}F_{X}(x)(1-F_{X}(x))\,\mathrm{d}x.$
A (law-invariant) signed Choquet integral, also called a distortion riskmetric, is a mapping of the form
$\rho_{h}:X\mapsto\int_{0}^{\infty}h(\mathbb{P}(X>x))\,\mathrm{d}x+\int_{-\infty}^{0}(h(\mathbb{P}(X>x))-h(1))\,\mathrm{d}x,$
where $h:[0,1]\to\mathbb{R}$ is a function of bounded variation satisfying $h(0)=0$, called a distortion function. The functional $\mathrm{GD}$ belongs to this class and has a distortion function $h:[0,1]\to\mathbb{R}$ given by h(t)=t(1t)h(t)=t(1-t).
The Gini coefficient $\mathrm{GC}:L^{1}_{+}\to[0,1]$ is defined as
$\mathrm{GC}(X)=\frac{\mathrm{GD}(X)}{\mathbb{E}[X]}=\frac{\int_{0}^{\infty}F_{X}(t)(1-F_{X}(x))\,\mathrm{d}x}{\int_{0}^{\infty}(1-F_{X}(x))\,\mathrm{d}x}.$
The range of $\mathrm{GC}$ is $[0,1)$. An alternative way of defining $\mathrm{GC}$ is through the Lorenz curve (Gastwirth (1971)),
$L_{F_{X}}(p)=\frac{\int_{0}^{p}F_{X}^{-1}(t)\,\mathrm{d}t}{\int_{0}^{1}F_{X}^{-1}(t)\,\mathrm{d}t},~~p\in[0,1],$
and GC is twice the area between $L_{F_{X}}$ and the identity on $[0,1]$,
$\mathrm{GC}(X)=2\int_{0}^{1}(p-L_{F_{X}}(p))\,\mathrm{d}p=1-2\int_{0}^{1}L_{F_{X}}(p)\,\mathrm{d}p.$
We generally refer to both GD and GC as Gini indices. Since GC and GD are connected by the simple relation $\mathrm{GC}=\mathrm{GD}/\mathbb{E}$, our axiomatic framework in the next section will mainly focus on properties satisfied by GD.
3 An axiomatic framework for higher-order Gini indices
3.1 Axioms of the Gini indices
Our aim is to generalize GD and GC to a family of flexible indices, while keeping the nice properties of GD and GC intact. For this purpose, we will propose several axioms satisfied by GD, and then look for functionals that also satisfy these axioms. We assume that the domain of the indices is $\mathcal{X}=L^{q}$ for some $q\geq 1$, and denote by $\mathcal{X}_{+}=L^{q}_{+}$.
The first observation, clearly from (2), is that GD can be written as the expected value of a function of two iid copies $X_{1},X_{2}$ from $X$. That is, there exists a function $f:\mathbb{R}^{2}\to\mathbb{R}$ such that
$\mathrm{GD}(X)=\mathbb{E}[f(X_{1},X_{2})].$
A natural generalizing of this property is to use $n\geq 3$ iid copies, which leads to the following axiom of Gini indices. We use $\rho:\mathcal{X}\to\mathbb{R}$ for a general functional as a candidate for a generalization of GD.
  1. [A1]
    Sample representability: There exists $f:\mathbb{R}^{n}\to\mathbb{R}$ for some $n\in\mathbb{N}$ such that $\rho(X)=\mathbb{E}[f(X_{1},\dots,X_{n})]$ for all $X\in\mathcal{X}$, where $X_{1},\dots,X_{n}$ are iid copies of $X$.
Besides [A1], GD satisfies many other natural axioms, which we list below.
  1. [A2]
    Symmetry: $\rho(X)=\rho(-X)$ for $X\in\mathcal{X}$.
  2. [A3]
    Comonotonic additivity: $\rho(X+Y)=\rho(X)+\rho(Y)$ if $X$ and $Y$ are comonotonic. [^3] Random variables $X$ and $Y$ are comonotonic if there exists $\Omega_{0}\in\mathcal{A}$ with $\mathbb{P}\left(\Omega_{0}\right)=1$ such that for all $\omega,\omega^{\prime}\in\Omega_{0}$, $\left(X(\omega)-X\left(\omega^{\prime}\right)\right)\left(Y(\omega)-Y\left(\omega^{\prime}\right)\right)\geq 0.$
  3. [A4]
    Uniform norm continuity: For any $\epsilon>0$, there exists $\delta>0$ such that $|\rho(X)-\rho(Y)|\leq\epsilon$ whenever $\operatorname*{ess\text{-}sup}|X-Y|\leq\delta$, where $\operatorname*{ess\text{-}sup}$ means the essential supremum.
  4. [A5]
    Nonnegativity: $\rho(X)\geq 0$ for $X\in\mathcal{X}$ and $\rho(X)=0$ if and only if $X$ is a constant.
  5. [A6]
    Location invariance : $\rho(X+c)=\rho(X)$ for all $c\in\mathbb{R}$ and $X\in\mathcal{X}$.
  6. [A7]
    Positive homogeneity: $\rho(\lambda X)=\lambda\rho(X)$ for all $\lambda\in(0,\infty)$ and $X\in\mathcal{X}$.
  7. [A8]
    Convexity: $\rho(\lambda X+(1-\lambda)Y)\leq\lambda\rho(X)+(1-\lambda)\rho(Y)$ for all $\lambda\in[0,1]$ and $X,Y\in\mathcal{X}$.
  8. [A9]
    Subadditivity: $\rho(X+Y)\leq\rho(X)+\rho(Y)$ for all $X,Y\in\mathcal{X}$.
  9. [A10]
    Convex-order consistency: $\rho(X)\leq\rho(Y)$ for all $X,Y\in\mathcal{X}$ whenever $X\leq_{\rm cx}Y$. [^4] A random variable $X$ is said to be smaller than a random variable $Y$ in convex order, denoted by $X\leq_{\rm cx}Y$, if $\mathbb{E}[\phi(X)]\leq\mathbb{E}[\phi(Y)]$ for all convex $\phi:\mathbb{R}\rightarrow\mathbb{R}$, provided that both expectations exist.
  10. [A11]
    Mixture-concavity: $\rho$ is concave as a mapping from $\mathcal{M}_{\rho}$ to $\mathbb{R}$, that is, $\rho(\lambda F+(1-\lambda)G)\geq\lambda\rho(F)+(1-\lambda)\rho(G)$ for all $F,G\in\mathcal{M}_{\rho}$ and $\lambda\in[0,1]$.
  11. [A12]
    Normalization: $\{\rho(X)/\mathbb{E}[X]:X\in\mathcal{X}_{+}\}=[0,1)$.
Axioms [A1]–[A12] are all satisfied by GD, and we will refer them to as the Gini axioms. Some axioms imply the others. For instance, positive homogeneity [A7] and convexity [A8] imply subadditivity [A9].
Note that [A12] is made so that $\rho/\mathbb{E}$ can be readily used as the corresponding relative index playing the role of GC. Therefore, the Gini axioms are made for the pair $(\rho,\rho/\mathbb{E})$ in place of $(\mathrm{GD},\mathrm{GC})$, although the mathematical statements for $\rho$ are simpler to present, except for [A12]. The economic interpretation of some of these axioms will be discussed in Section 4.
3.2 Axiomatic characterization
Our aim is to generalize the Gini deviation and the Gini coefficient to higher order. In what follows, always assume $n\geq 2$. To establish a representation with as few axioms as possible, we will only assume [A1]–[A4] in the main direction of the result.
Theorem 1.
For a mapping $\rho:\mathcal{X}\to\mathbb{R}$, it satisfies [A1]–[A4] if and only if there exist an integer $n\in\mathbb{N}$ and $(a_{1},\dots,a_{n})\in\mathbb{R}^{n}$ such that
$\rho(X)=\sum_{i=1}^{n}a_{i}{\rm GD}_{i}(X),~~X\in\mathcal{X},$ (3)
where $\mathrm{GD}_{1}(X)=\mathrm{GD}(X)$, and for $i=2,3,\dots,n$,
$\mathrm{GD}_{i}(X)=\frac{1}{i}\mathbb{E}\left[\max\{X_{1},\dots,X_{i}\}-\min\{X_{1},\dots,X_{i}\}\right],$ (4)
with $X_{1},\dots,X_{n}$ being iid copies of $X$. Moreover, if $(a_{1},\dots,a_{n})\in\Delta_{n}$, then $\rho(X)$ in (3) satisfies all of the Gini axioms [A1]–[A12].
The most important part of the proof of Theorem 1 is the necessity of (3). This involves two steps. The first step establishes a representation via signed Choquet integrals. This follows from results in Wang et al. (2020a), which uses [A3], [A4], and law invariance (implied by [A1]) to establish the representation. The second step identifies symmetric ([A2]) signed Choquet integrals that can be represented by expected values of functions of iid observations ([A1]), a new property to the literature that requires sophisticated analysis.
Inspired by the representation (3) in Theorem 1, we define the $n$-th order Gini deviation for $n\geq 2$ as the functional $\mathrm{GD}_{n}$ in (4), and without loss of generality its domain is chosen as $L^{1}$. That is,
$\mathrm{GD}_{n}(X)=\frac{1}{n}\mathbb{E}\left[\max\{X_{1},\dots,X_{n}\}-\min\{X_{1},\dots,X_{n}\}\right],~~~~X\in L^{1}.$
Note that (2) can be rewritten as
$\mathrm{GD}(X)=\frac{1}{2}\mathbb{E}\left[\max\{X,X^{\prime}\}-\min\{X,X^{\prime}\}\right],$
and therefore, $\mathrm{GD}_{2}=\mathrm{GD}$. Following this, the $n$-th order Gini coefficient $\mathrm{GC}_{n}$ for $n\geq 2$, is defined by, as in Gavilan-Ruiz et al. (2024),
$\mathrm{GC}_{n}(X)=\frac{\mathrm{GD}_{n}(X)}{\mathbb{E}[X]},~~~~X\in L^{1}_{+}.$
Note that $\mathrm{GC}_{n}$ has a natural range of $[0,1)$ due to [A12], and this fact has been observed by Gavilan-Ruiz et al. (2024). The following example explains the extreme distributions, and also justifies that $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ bear the same interpretation as $\mathrm{GD}$ and $\mathrm{GC}$ when it comes to measuring wealth or income inequality, as the worst-case scenario is attained asymptotically by the distribution in which a tiny ($\epsilon$ in the example) proportion of the population has all the wealth or income.
Example 1.
Suppose that $X_{\epsilon}$ follows a Bernoulli distribution with mean $\epsilon>0$. From the definition of $\mathrm{GD}_{n}$, we can compute
$\mathrm{GC}_{n}(X_{\epsilon})=\frac{\mathrm{GD}_{n}(X_{\epsilon})}{\mathbb{E}[X_{\epsilon}]}=\frac{1}{\epsilon}\int_{1-\epsilon}^{1}(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t=\frac{1}{n\epsilon}\left(1-\epsilon^{n}-(1-\epsilon)^{n}\right),$
and by sending $\epsilon\downarrow 0$, we get
$\mathrm{GC}_{n}(X_{\epsilon})=\frac{1}{n\epsilon}\left(n\epsilon+O(\epsilon^{2})\right)\to 1.$
Some members in the family of high-order Gini deviations coincide. In addition to $\mathrm{GD}_{2}=\mathrm{GD}$, we can also check $\mathrm{GD}_{2}=\mathrm{GD}_{3}$ by the equality
$\max\{X_{1},X_{2},X_{3}\}-\min\{X_{1},X_{2},X_{3}\}=\frac{1}{2}(|X_{1}-X_{2}|+|X_{1}-X_{3}|+|X_{2}-X_{3}|),$
which implies
$\mathrm{GD}_{3}(X)=\frac{1}{6}\mathbb{E}[|X_{1}-X_{2}|+|X_{1}-X_{3}|+|X_{2}-X_{3}|]=\frac{1}{2}\mathbb{E}[|X_{1}-X_{2}|]=\mathrm{GD}_{2}(X).$
Nevertheless, $\mathrm{GD}_{4}\neq\mathrm{GD}_{3}$ in general, and $\mathrm{GD}_{n}$ decreases in $n$ (see Appendix C).
Inspired by the axiomatic characterization of indices satisfying the Gini axioms in Theorem 1, we focus on the families $(\mathrm{GD}_{n})_{n\geq 2}$ and $(\mathrm{GC}_{n})_{n\geq 2}$ of higher-order Gini indices in the rest of the paper. In particular, $\mathrm{GC}_{n}$ for some different choices of $n\geq 2$ will be our main tool for measuring inequality as a complement to the classic $\mathrm{GC}$, while we should keep in mind that their convex combinations also satisfy the Gini axioms and can be used for the same purpose.
Remark 1.
In Theorem 1, we have shown that if $(a_{1},\dots,a_{n})$ is in $\Delta_{n}$, then all the Gini axioms are satisfied. One may wonder whether the reverse holds true, that is, whether for $\rho$ in (3), [A1]–[A12] jointly imply $(a_{1},\dots,a_{n})\in\Delta_{n}$. It turns out that this is not true, even removing the redundant terms (e.g., $\mathrm{GD}_{1}=\mathrm{GD}_{2}=\mathrm{GD}_{3}$). For an example, define $\rho(X)=2{\rm GD}_{2}(X)-{\rm GD}_{4}(X)$. By Proposition 1, we know that $\rho$ is a signed Choquet integral whose distortion function is given by
$h(t)=(1-t^{2}-(1-t)^{2})-\frac{1}{4}(1-t^{4}-(1-t)^{4}).$
One can check that $h$ is concave on $[0,1]$, and thus, $\rho$ is convex. Using this fact, we can check all of [A5]–[A12] by standard computations.
4 Properties of the higher-order Gini indices
4.1 Economic properties of higher-order GC
The next theorem gives some useful and economically interpretable properties of $\mathrm{GC}_{n}$, similar to the properties of $\mathrm{GD}_{n}$ in Theorem 1, but with different forms.
Theorem 2.
The mapping $\mathrm{GC}_{n}:L^{1}_{+}\to\mathbb{R}$ satisfies
  1. (i)
    Scale invariance: $\mathrm{GC}_{n}(\lambda X)=\mathrm{GC}_{n}(X)$ for all $\lambda\in(0,\infty)$ and $X\in L^{1}_{+}$;
  2. (ii)
    Convex-order consistency: $\mathrm{GC}_{n}(X)\leq\mathrm{GC}_{n}(Y)$ for all $X,Y\in L^{1}_{+}$ whenever $X\leq_{\rm cx}Y$;
  3. (iii)
    Quasi-convexity: $\mathrm{GC}_{n}(\lambda X+(1-\lambda)Y)\leq\max\{\mathrm{GC}_{n}(X),\mathrm{GC}_{n}(Y)\}$ for all $\lambda\in[0,1]$ and $X,Y\in L^{1}_{+}$;
  4. (iv)
    Mixture-quasi-concavity: $\mathrm{GC}_{n}$ is quasi-concave as a mapping from $\mathcal{M}^{1}_{+}$ to $\mathbb{R}$, that is, $\mathrm{GC}_{n}(\lambda F+(1-\lambda)G)\geq\min\{\mathrm{GC}_{n}(F),\mathrm{GC}_{n}(G)\}$ for all $F,G\in\mathcal{M}^{1}_{+}$ and $\lambda\in[0,1]$.
The four properties of $\mathrm{GC}_{n}$ in Theorem 2 have natural economic interpretations as measures of inequality, similarly to the case of $\mathrm{GC}$. To explain these interpretations, we use measuring income inequality as the main example.
  1. 1.
    Scale invariance means that a rescaling of the income distribution, e.g., multiplying by a currency exchange rate, does not affect the measurement. Therefore, $\mathrm{GC}_{n}$ can be evaluated in either the local currency or a standard one, such as the US dollars.
  2. 2.
    Convex-order consistency of $\mathrm{GC}_{n}$ means that a more spread-out distribution has a larger $\mathrm{GC}_{n}$, confirming the role of $\mathrm{GC}_{n}$ as a measure of distributional dispersion. In particular, if the income process (with respect to time) in a country forms a martingale (i.e., the expected income in the next year of an individual is equal to their income this year), then $\mathrm{GC}_{n}$ will increase in time.
  3. 3.
    Quasi-convexity of $\mathrm{GC}_{n}$ means that if we aggregate income $X$ and income $Y$ from two populations of equal size in pair, for instance, forming households from pairs of individuals, then the overall value of $\mathrm{GC}_{n}$ is smaller than the maximum of $\mathrm{GC}_{n}(X)$ and $\mathrm{GC}_{n}(Y)$. This means that aggregating groups of individuals with similar level of inequality generally reduces inequality.
  4. 4.
    Mixture-quasi-concavity of $\mathrm{GC}_{n}$ means that if we merge two populations with different income distributions $X$ and $Y$ to make a bigger population (for instance, forming a larger region from two small regions), then the overall value of $\mathrm{GC}_{n}$ is larger than the minimum of $\mathrm{GC}_{n}(X)$ and $\mathrm{GC}_{n}(Y)$. This suggests that a larger country likely has a larger $\mathrm{GC}_{n}$. This is a natural condition for measuring inequality. For example, two populations that are perfectly equal among themselves (e.g., one rich group and one poor group) can exhibit significant inequality when assessed as a combined population.
Convex-order consistency is also used to define strong risk aversion, that is, aversion to mean-preserving spreads in the sense of Rothschild and Stiglitz (1970). Strong risk aversion of preference models is widely studied in decision theory. For instance, Schmidt and Zank (2008) provided a characterization of strong risk aversion within the framework of cumulative prospect theory.
The above point 3 (quasi-convexity) and point 4 (mixture-quasi-concavity) should not be seen as conflicting, as summing two income random variables and merging two populations are very different operations and have opposite effects. Indeed, there is a fundamental conflict between convexity and mixture-convexity (see Wang and Wu (2025, Proposition 6)), whereas convexity and mixture-concavity are equivalent for Signed Choquet integrals (Wang et al. (2020b, Theorem 3)).
4.2 Representation as signed Choquet integrals
In the next proposition, we represent $\mathrm{GD}_{n}$ as a signed Choquet integral. This is useful in proving several properties of $\mathrm{GD}_{n}$ in Theorem 1. In what follows, let $U_{X}$ be a uniform random variable such that $F_{X}^{-1}(U_{X})=X$ almost surely, and its existence is guaranteed by Föllmer and Schied (2016, Lemma A.32).
Proposition 1.
The mapping $\mathrm{GD}_{n}:L^{1}\to\mathbb{R}$ satisfies the following properties.
  1. (i)
    It has the quantile representation
    $\mathrm{GD}_{n}(X)=\int_{0}^{1}F_{X}^{-1}(t)(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t,~~~~X\in L^{1}.$ (5)
  2. (ii)
    It is a signed Choquet integral with a concave distortion function
    $h_{n}(t)=\frac{1}{n}\left(1-t^{n}-(1-t)^{n}\right),~~~~t\in[0,1].$ (6)
Figure 4 depicts $h_{n}(t)$ in (6) and its derivative $h_{n}^{\prime}(t)$ for various values of $n$. As $n$ increases, $h_{n}(t)$ becomes flatter over the interval $[0,1]$, exhibiting increasingly sharp transitions near the endpoints $t=0$ and $t=1$. Note that $h_{n}^{\prime}(t)$ serves as the “probability weight” assigned to $t$ in the calculation of the Choquet expectation; see, e.g. Quiggin (1982) and Gilboa and Schmeidler (1989). From this perspective, $\mathrm{GD}_{n}$ also reflects an increasing sensitivity to tail behavior and concentration effects.
Figure 4: The distortion function $h_{n}(t)=\left(1-t^{n}-(1-t)^{n}\right)/n$ (left panel) and its derivative $h_{n}^{\prime}(t)=(1-t)^{n-1}-t^{n-1}$ (right panel) for $n=2,5,10,20,50$.
Remark 2.
Let $\rho_{n}$ be the power distortion risk measure defined by
$\rho_{n}(X)=\int_{0}^{1}nt^{n-1}F^{-1}_{X}(t)\,\mathrm{d}t,~~~X\in L^{1}.$
It is well known that $\rho_{n}$ is a coherent risk measure (Artzner et al. (1999)). The functional $\mathrm{GD}_{n}$ is connected to $\rho_{n}$ via
$\mathrm{GD}_{n}(X)=\frac{\rho_{n}(X)+\rho_{n}(-X)}{n}.$
Remark 3.
From (5), we can easily generalize the definition of $\mathrm{GD}_{n}$ to real numbers $n>1$, and these indices are also signed Choquet integrals. Similarly, one can define $\mathrm{GC}_{n}$ for all real numbers $n>1$. Nevertheless, for a non-integer $n$, sample representability [A1] is no longer satisfied by $\mathrm{GD}_{n}$ in (5). For this reason, we focus integer values of $n$.
4.3 Comparing Gini deviations and the standard deviation
The inequality $0\leq\mathrm{GD}(X)/\mathrm{SD}(X)\leq 3^{-1/2}$ for all nonconstant $X\in L^{2}$ is known as Glasser’s inequality (Glasser (1962)), where $\mathrm{SD}(X)$ denotes the standard deviation of $X$. For the general $\mathrm{GD}_{n}$, we establish the following bounds. The proof follows a similar approach to Theorem 5 of Pesenti et al. (2025).
Proposition 2.
For $n\geq 2$ and nonconstant $X\in L^{2}$, we have
$0\leq\frac{\mathrm{GD}_{n}(X)}{\mathrm{SD}(X)}\leq\sqrt{\frac{2}{2n-1}-\frac{2((n-1)!)^{2}}{(2n-1)!}},$ (7)
and these bounds are sharp.
Note that the right-hand side of (7) decreases in $n$ and it is very close to $n^{-1/2}$ for $n$ large. Its maximum value is $3^{-1/2}$ when $n=2,3$.
Next, we consider the bounds on $\mathrm{GD}_{n}(X)/\mathrm{GD}_{m}(X)$. In fact, we can derive a more general result for bounds on the ratio of any two deviation measures defined via the Choquet integrals. Let
[…]
For $h,g\in\mathcal{H}$ and a nonconstant $X$ such that $\rho_{h}(X)$ and $\rho_{g}(X)$ are finite, we have
$\frac{\rho_{h}(X)}{\rho_{g}(X)}=\frac{\int_{-\infty}^{\infty}h(\mathbb{P}(X>x))\,\mathrm{d}x}{\int_{-\infty}^{\infty}g(\mathbb{P}(X>x))\,\mathrm{d}x}\in\left[\inf_{t\in(0,1)}\frac{h(t)}{g(t)},\sup_{t\in(0,1)}\frac{h(t)}{g(t)}\right].$ (8)
The proof of (8) follows directly from the definition of Choquet integrals. Moreover, we can see that these bounds are sharp as the value $h(t)/g(t)$ for $t\in(0,1)$ is attainable by a Bernoulli random variable with mean $t$. A specialization of (8) to $\mathrm{GD}_{n}$ leads to the following bounds.
Proposition 3.
Let $L^{1}_{*}$ be the set of nonconstant elements of $L^{1}$. For $2\leq m\leq n$, we have
$\min_{X\in L^{1}_{*}}\frac{\mathrm{GD}_{n}(X)}{\mathrm{GD}_{m}(X)}=\frac{m(1-2^{1-n})}{n(1-2^{1-m})};~~~~~\sup_{X\in L^{1}_{*}}\frac{\mathrm{GD}_{n}(X)}{\mathrm{GD}_{m}(X)}=1.$ (9)
The same bounds in Proposition 3 also apply to $\mathrm{GC}_{n}(X)/\mathrm{GC}_{m}(X)$ for $X\in L^{1}_{+}$. In particular, these bounds imply $1\leq\mathrm{GC}_{3}(X)/\mathrm{GC}_{2}(X)\leq 1$, thus $\mathrm{GC}_{3}=\mathrm{GC}_{2}$, and $7/8\leq\mathrm{GC}_{4}(X)/\mathrm{GC}_{3}(X)\leq 1$.
5 Elicitability and multi-observation elicitability
Statistical elicitability is an important property for statistical functionals, popular in forecasting, risk management, and machine learning; see Gneiting (2011) for a thorough study. It refers to the existence of a scoring function for the forecasted value of a risk functional and realized value of future observations, so that the mean of the scoring function attains its minimum value if and only if the value of the risk functional is truly forecasted. Comparative backtests, for which elicitability is a necessary condition, are discussed by Nolde and Ziegel (2020) as an alternative to the traditional backtests.
In this section, we mainly treat $\rho$ as mappings from $\mathcal{M}^{q}$ to $\mathbb{R}$ for some $q\geq 1$, instead of $L^{p}$ to $\mathbb{R}$ for some $q\geq 1$. The notation $f:A\rightrightarrows B$” means that $f$ maps each element of $A$ to a subset of $B$. Moreover, we identify a singleton in $\mathbb{R}^{k}$ with its element. For instance, if $\rho:\mathcal{M}\rightrightarrows\mathbb{R}^{k}$ takes value only in singletons, then we treat $\rho:\mathcal{M}\to\mathbb{R}^{k}$.
Definition 1.
For $\mathcal{M}\subseteq\mathcal{M}_{\rho}$, the mapping $\rho:\mathcal{M}_{\rho}\rightrightarrows\mathbb{R}^{k}$ is $\mathcal{M}$-elicitable if there exists a function $S:\mathbb{R}^{k+1}\rightarrow\mathbb{R}$ such that for all $F\in\mathcal{M}$,
$\rho(F)=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{k}}\int_{-\infty}^{\infty}S(\mathbf{x},y)\mathrm{d}F(y).$
We omit $\mathcal{M}$” in $\mathcal{M}$-elicitability” when $\mathcal{M}$ is the domain $\mathcal{M}_{\rho}$ of the mapping $\rho$. The function $S$ is called a score function.
It is easy to check that none of GD or GC is elicitable. In fact, as shown by Wang and Wei (2020), all deviation measures (including GD) are not elicitable.
Inspired by the definition of GD, $\mathrm{GD}_{n}$ is connected to another version of elicitability. This idea of multi-observation elicitability appeared first in Casalaina-Martin et al. (2017) and it is further studied by Frongillo et al. (2017).
Definition 2.
For $\mathcal{M}\subseteq\mathcal{M}_{\rho}$, a mapping $\rho:\mathcal{M}_{\rho}\rightrightarrows\mathbb{R}^{k}$ is $n$-observation $\mathcal{M}$-elicitable if there exists a function $S:\mathbb{R}^{k+n}\rightarrow\mathbb{R}$ such that
$\rho(F)=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{k}}\int_{\mathbb{R}^{n}}S(\mathbf{x},\mathbf{y})\mathrm{d}F^{n}(\mathbf{y}),\quad F\in\mathcal{M},$
where $F^{n}\left(y_{1},\ldots,y_{n}\right)=\prod_{j=1}^{n}F\left(y_{j}\right)$.
Two prominent examples of 2-observation elicitable functionals are the variance (Example 2 below) and the Gini deviation (Theorem 3 below).
Example 2.
Similar to the Gini deviation in (2), the variance functional can be written as
$\mathrm{Var}(X)=\frac{1}{2}\mathbb{E}[(X-X^{\prime})^{2}],$
where $X^{\prime}$ is an iid copy of $X$. The variance is $2$-observation $\mathcal{M}^{4}$-elicitable with the score function (e.g., Casalaina-Martin et al. (2017))
$S(x,y_{1},y_{2})=(2x-(y_{1}-y_{2})^{2})^{2}.$
This can be checked directly from
$\operatorname*{arg\,min}_{x\in\mathbb{R}}\mathbb{E}\left[(2x-(X-X^{\prime})^{2})^{2}\right]=\frac{1}{2}\mathbb{E}[(X-X^{\prime})^{2}]=\mathrm{Var}(X).$
Clearly, $1$-observation elicitability coincides with elicitability in Definition 1. Moreover, the property of $n$-observation elicitability gets weaker as $n$ grows.
Proposition 4.
If $m\geq n$, then $n$-observation elicitability implies $m$-observation elicitability.
This proposition follows directly by noticing that the function $S_{m}$ given by $S_{m}(x,y_{1},\dots,y_{m})=S_{n}(x,y_{1},\dots,y_{n})$, where $S_{n}$ is the $n$-observation score function, is indeed an $m$-observation score function.
Next, we establish two score functions for $\mathrm{GD}_{n}$.
Theorem 3.
The mapping $\mathrm{GD}_{n}$ is $n$-observation $\mathcal{M}^{2}$-elicitable with the score function
$S\left(x,y_{1},\dots,y_{n}\right)=\left(nx-\max\{y_{1},\dots,y_{n}\}+\min\{y_{1},\dots,y_{n}\}\right)^{2}.$
Moreover, $\mathrm{GD}_{n}$ is $n$-observation $\mathcal{M}^{1}$-elicitable with the score function
$S^{*}\left(x,y_{1},\dots,y_{n}\right)=nx^{2}-2x\left(\max\{y_{1},\dots,y_{n}\}-\min\{y_{1},\dots,y_{n}\}\right).$
In particular, $\mathrm{GD}_{2}$ is $2$-observation $\mathcal{M}^{2}$-elicitable by $(x,y_{1},y_{2})\mapsto(2x-|y_{1}-y_{2}|)^{2}$ and $2$-observation $\mathcal{M}^{1}$-elicitable by $(x,y_{1},y_{2})\mapsto x^{2}-x|y_{1}-y_{2}|$.
Similarly to the classic notion of elicitability, the score function for an $n$-observation elicitable functional is not unique in general, as suggested by Theorem 3.
The next result shows $n$-observation elicitability of signed Choquet integrals with a polynomial distortion function.
Theorem 4.
A signed Choquet integral, with distortion function $h$ that is a polynomial of degree $n$ or less, is $n$-observation $\mathcal{M}^{1}$-elicitable.
When $n$ is an odd number, we can check that the $n$-th degree term $t^{n}$ in the distortion function $h_{n}$ in (6) get canceled. Therefore, in this case, by using Theorem 4, $\mathrm{GD}_{n}$ is (n1)(n-1)-observation elicitable, which is a stronger condition than $n$-observation elicitability, as shown in Proposition 4. Note that the score function for n1n-1 observations is different from the one obtained in Theorem 3 for $n$ observations.
The functional $\mathrm{GC}_{n}$ is also $n$-observation elicitable, as the ratio of expectations, following the same idea as in Theorem 3.2 of Gneiting (2011).
Proposition 5.
The mapping $\mathrm{GC}_{n}$ is $n$-observation $\mathcal{M}^{1}$-elicitable with the score function
$S\left(x,y_{1},\dots,y_{n}\right)=x^{2}y_{1}-\frac{2x}{n}(\max\{y_{1},\dots,y_{n}\}-\min\{y_{1},\dots,y_{n}\}).$
In particular, $\mathrm{GC}_{2}$ is $2$-observation $\mathcal{M}^{1}$-elicitable by $S(x,y_{1},y_{2})=x^{2}y_{1}-x|y_{1}-y_{2}|$.
The elicitability results allow for obtaining $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ through empirical risk minimization. In the next section, we will study estimators of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ through their Choquet integral representation.
6 Estimation and inference
In this section, we study empirical estimators of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ and establish their consistency and asymptotic normality, which can be used for point estimation and hypothesis testing.
Let $X_{1},\ldots,X_{N}$ be a sample of independent observations, and denote by $X_{(1)}\leq\cdots\leq X_{(N)}$ their order statistics. The empirical distribution function is given by $\widehat{F}_{N}(x)=\frac{1}{N}\sum_{j=1}^{N}\mathbf{1}_{\{X_{j}\leq x\}},$ and its associated quantile function satisfies $\widehat{F}_{N}^{-1}(i/N)=X_{(i)}$ for $i=1,\ldots,N$. As a standard approach for estimating L-statistics and distortion risk measures (e.g., Huber and Ronchetti (2009) and McNeil et al. (2015)), the empirical estimator of $\mathrm{GD}_{n}(X)$ is then defined as
$\widehat{\mathrm{GD}}_{n}(N)=\frac{1}{N}\sum_{i=1}^{N}X_{(i)}\left(\left(\frac{i}{N}\right)^{n-1}-\left(1-\frac{i}{N}\right)^{n-1}\right),$
while the empirical estimator of $\mathrm{GC}_{n}(X)$ takes the form
$\widehat{\mathrm{GC}}_{n}(N)=\frac{1}{\overline{X}}\cdot\widehat{\mathrm{GD}}_{n}(N),\quad\text{where }\overline{X}=\frac{1}{N}\sum_{i=1}^{N}X_{i}.$
We make following standard regularity assumption on the distribution of the random variable $X$.
Assumption 1.
The distribution $F$ of $X\in L^{1}$ has a support that is a convex set and has a positive density function $f$ on the support. Denote by $\tilde{f}=f\circ F^{-1}$.
Theorem 5.
Suppose that $X_{1},\dots,X_{N}\in L^{1}_{+}$ are an iid sample from $X\in L^{1}_{+}$ and Assumption 1 holds. Then, we have $\widehat{\mathrm{GD}}_{n}(N)\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\mathrm{GD}_{n}(X)$ and $\widehat{\mathrm{GC}}_{n}(N)\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\mathrm{GC}_{n}(X)$ as $n\to\infty.$ Moreover, if $X\in L^{\gamma}_{+}$ for some $\gamma>2$, then we have
$\sqrt{N}\left(\widehat{\mathrm{GD}}_{n}(N)-\mathrm{GD}_{n}(X)\right)$ $\stackrel{{\scriptstyle\mathrm{d}}}{{\rightarrow}}\mathrm{N}\left(0,\sigma^{2}_{\mathrm{GD}_{n}}\right);$
$\sqrt{N}\left(\widehat{\mathrm{GC}}_{n}(N)-\mathrm{GC}_{n}(X)\right)$ $\stackrel{{\scriptstyle\mathrm{d}}}{{\rightarrow}}\mathrm{N}\left(0,\sigma^{2}_{\mathrm{GC}_{n}}\right),$
where
$\sigma^{2}_{\mathrm{GD}_{n}}$ $=\int_{0}^{1}\int_{0}^{1}\frac{\phi_{n}(s)\phi_{n}(t)(s\wedge t-st)}{\tilde{f}(s)\tilde{f}(t)}\mathrm{d}t\mathrm{d}s,$ (10)
$\sigma^{2}_{\mathrm{GC}_{n}}$ $=\int_{0}^{1}\int_{0}^{1}\frac{(\phi_{n}(s)-\mathrm{GC}_{n}(X))(\phi_{n}(t)-\mathrm{GC}_{n}(X))(s\wedge t-st)}{(\mathbb{E}[X])^{2}\tilde{f}(s)\tilde{f}(t)}\mathrm{d}t\mathrm{d}s,$ (11)
and $\phi_{n}(s)=s^{n-1}-(1-s)^{n-1}$ for $s\in[0,1]$.
The computational complexity of the empirical estimator $\widehat{\mathrm{GD}}_{n}(N)$ is dominated by the sorting step required to compute the empirical distribution function. Sorting the sample $X_{1},X_{2},\dots,X_{N}$ takes $O(N\log N)$ time using efficient sorting algorithms such as quicksort or mergesort. After sorting, the calculation of the empirical distribution and the estimator itself requires only $O(N)$ operations. Therefore, the overall time complexity of $\widehat{\mathrm{GD}}_{n}(N)$ is $O(N\log N)$.
In what follows, we present some simulation results based on Theorem 5. Simulation results are presented in the case of log-normal distribution $\mathrm{LN}(0,1)$ and $\text{Pareto}(3,2)$ (with tail index $3$). Let the sample size $N=5000$, and we repeat the procedure 2000 times. We assume $n=5$.
Figure 5: Values of $\widehat{\mathrm{GD}}_{5}(N)$ (left panel) and $\widehat{\mathrm{GC}}_{5}(N)$ (right panel) for $\mathrm{LN}(0,1)$
Figure 6: Values of $\widehat{\mathrm{GD}}_{5}(N)$ (left panel) and $\widehat{\mathrm{GC}}_{5}(N)$ (right panel) for $\mathrm{Pareto}(3,2)$
In Figure 5, the sample is drawn from the log-normal distribution $\mathrm{LN}(0,1)$. We observe that the empirical estimates of $\mathrm{GD}_{5}(X)$ align quite well with the density of $\mathrm{N}(0.74,2.33/N)$, while the empirical estimates of $\mathrm{GC}_{5}(X)$ match the density of $\mathrm{N}(0.45,0.19/N)$. Notably, the asymptotic variance of $\mathrm{GC}_{5}(X)$ is smaller than that of $\mathrm{GD}_{5}(X)$. In Figure 6, the sample is drawn from the Pareto distribution $\mathrm{Pareto}(3,2)$. We observe that the empirical estimates of $\mathrm{GD}_{5}(X)$ closely match the density of $\mathrm{N}(0.52,1.83/N)$, and the empirical estimates of $\mathrm{GC}_{5}(X)$ align with the density of $\mathrm{N}(0.17,0.13/N)$.
Figure 7: Asymptotic variance for $\mathrm{GD}_{n}(X)$ (left panel) and $\mathrm{GC}_{n}(X)$ (right panel) for $X\sim\mathrm{LN}(0,1)$
Figure 8: Asymptotic variance for $\mathrm{GD}_{n}(X)$ (left panel) and $\mathrm{GC}_{n}(X)$ (right panel) for $X\sim\mathrm{Pareto}(3,2)$
In addition, we illustrate how the value of asymptotic variance varies as $n$ ranges from 2 to 20. As shown in Figures 7 and 8, the asymptotic variance of $\mathrm{GD}_{n}(X)$ decreases as $n$ increases. However, the asymptotic variance of $\mathrm{GC}_{n}(X)$ does not necessarily exhibit such monotonic behavior. For example, when $X\sim\mathrm{LN}(0,1)$, the asymptotic variance of $\mathrm{GC}_{n}(X)$ initially increases and then decreases.
7 Empirical analysis
In this section, we present an empirical analysis by computing high-order Gini indices, $\mathrm{GC}_{n}$, for both wealth and income across countries and regions over time, and by comparing these results with the conventional case of $n=2$. As suggested by Theorem 1, one could also consider linear combinations of $\mathrm{GC}_{n}$, but this will lead to similar results, which we omit.
Our analysis is based on data from the World Inequality Database (WID), with all monetary figures converted to United States dollars to ensure cross-country comparability.
Figure 9: Values of $\mathrm{GC}_{n}$ based on household wealth for the United States, China, India, German, Brazil and the South Africa from 2000 to 2023, for $n=2$, 5, 10, and 20.
Figure 10: Wealth shares held by the top 1% (left) and top 10% (right) in the United States, China, India, German, Brazil and the South Africa from 2000 to 2023.
We begin by examining the values of $\mathrm{GC}_{n}$ for six major economies—the US, China (mainland), India, German, Brazil and South Africa—over the period 2000 to 2023. The plots for Canada and the UK have already been presented in the Introduction and are therefore omitted here for clarity. Figure 9 reveals several notable patterns, similar to those observed in Firgure 3. For $n=2$, the values of $\mathrm{GC}_{2}$ for China, India, and German are quite close after 2010, indicating broadly similar levels of overall wealth inequality. In contrast, the South Africa and US consistently displays a much higher $\mathrm{GC}_{2}$, reflecting its well-documented concentration of wealth. The value for Brazil lies between these two groups, remaining below that of the US. As $n$ increases to 5, 10, and 20, the values of Brazil and the US become increasingly similar. Additionally, China’s $\mathrm{GC}_{n}$ values rise sharply, surpassing those of India and German and gradually approaching those of the United States. While the trajectories for India and Germany remain closely aligned across all values of $n$, a change emerges after 2010: the valuse of $\mathrm{GC}_{10}$ and $\mathrm{GC}_{20}$ for India slightly exceed those of Germany, reversing the pattern seen for $\mathrm{GC}_{2}$ and $\mathrm{GC}_{5}$.
To better interpret these results, we plot in Figure 10 the wealth shares held by the top 1% and top 10% in each country. South Africa maintains both high top-share levels and persistently elevated $\mathrm{GC}_{n}$ values, reaffirming its status as a highly unequal economy. China’s top 10% share exceeds those of India and Germany, which is consistent with the sharp rise in line with the pronounced increase in its $\mathrm{GC}_{n}$ values as $n$ increases. We also observe that the top 1% and top 10% wealth shares in Brazil are quite close to those in the US, helping to explain the behavior of its $\mathrm{GC}_{n}$: as $n$ increases, Brazil’s index rises significantly and the difference between Brazil and US almost vanishes. Furthermore, India’s top 1% and top 10% shares surpass those of Germany after 2004, which explains the shift in their relative $\mathrm{GC}_{n}$ patterns—while Germany initially has higher values at smaller $n$, India overtakes it as $n$ grows.
Figure 11: Values of $\mathrm{GC}_{n}$ based on post-tax national income for Africa, Asia, Europe, North America, South America, and Oceania from 2000 to 2023, for $n=2$, 5, 10, and 20.
Figure 12: Income shares held by the top 1% (left) and top 10% (right) in Africa, Asia, Europe, North America, South America and Oceania from 2000 to 2023.
Traditionally, GC is used for measuring income inequality. We present the values of $\mathrm{GC}_{n}$ for post-tax national income across six major world continents—Africa, Asia, Europe, North America, South America, and Oceania—over the period 2000 to 2023, with $n\in\{2,5,10,20\}$. [^5] WID provides the full income distribution for each continent, but not for each country. By Figure 11, we can see that for all $n\geq 2$, $\mathrm{GC}_{n}$ for Africa, Asia, and South America consistently exceed those of North America, Europe, and Oceania. While $\mathrm{GC}_{2}$ shows only minor differences among North America, Europe, and Oceania, the gap between North America and Oceania widens progressively with higher $n$, reflecting greater top-end income concentration consistent with its relatively large shares held by the top 1% and 10% income groups as shown in Figure 12. Further, from the perspective of $\mathrm{GC}_{2}$, Asia’s inequality level is close to South America’s and lower than Africa’s. However, when $n\geq 5$, Asia’s $\mathrm{GC}_{n}$ falls below those of both Africa and South America. South America’s $\mathrm{GC}_{n}$ values at higher $n$ surpass those of the other two regions, marking it as the region with the highest concentration of top incomes. This aligns with evidence showing that South America’s top 10% income shares exceed those of both Asia and Africa in Figure 12.
These findings highlight the key advantage of $\mathrm{GC}_{n}$: it uncovers differences that the traditional GC cannot detect. When overall inequality appears similar, higher-order $\mathrm{GC}_{n}$ values reveal significant disparities in top-tail income concentration. Thus, $\mathrm{GC}_{n}$ provides a more nuanced tool for assessing income concentration and its evolution over time, offering particular value to policymakers monitoring the impacts of tax reforms or addressing the accumulation of wealth among ultra-high-net-worth individuals.
From the empirical results, if one needs to pick $\mathrm{GC}_{n}$ for a fixed value of $n$ in complement to the standard $\mathrm{GC}_{2}$, then we generally recommend using $\mathrm{GC}_{n}$ with $n=10$, as using a small value of $n$ does not provide much different information than $\mathrm{GC}_{2}$, and using $n>10$ leads to qualitatively similar results. Nevertheless, in practice it may be helpful to compute $\mathrm{GC}_{n}$ for several different $n$, or an average of them, to get a more comprehensive picture on the inequality to be measured.
8 Conclusion
We provided an axiomatic study for a generalization of the classical Gini deviation and Gini coefficientas tools for inequality and dispersion measurement. We considered an axiom involving expectations of functions of $n\geq 2$ independent and identically distributed copies. By imposing the properties of symmetry, comonotonic additivity, and continuity, we obtained an axiomatic characterization showing that any such measure can be represented as an affine combination of the higher-order Gini deviations. By transitioning from pairwise to groupwise comparison structures, the generalized indices capture aspects of tail concentration and extremal behavior, different from the traditional two-observation formulation.
We analyze the economic properties of the higher-order Gini coefficient, and show that higher-order Gini deviations admit Choquet integral representations, satisfying key properties of coherent deviation and risk measures. The relationship between the high order Gini deviation and the standard deviation is explored. Furthermore, we establish their $n$-observation elicitability, bridging the gap between descriptive indices and decision-theoretic evaluation. Empirical analyses using global income and wealth data illustrate the strength of this framework: while the classical GC often masks nuanced disparities across regions or subpopulations, higher-order indices reveal meaningful differences—particularly among top earners. These findings highlight the value of $\mathrm{GC}_{n}$ as a diagnostic tool for policymakers, regulators, and data scientists facing increasingly complex distributional challenges.
As data environments grow more intricate, the flexibility of higher-order Gini indices make them promising instruments for capturing refined patterns of inequality, concentration, and systemic risk. In risk management, these indices provide coherent, tail-sensitive alternatives to variance-based risk measures. Their emphasis on upper-tail risks aligns naturally with the objectives of solvency analysis, reinsurance pricing, and regulatory capital assessment. Meanwhile, in machine learning, the multi-observation structure supports the design of loss functions that better capture group-level dispersion, promoting robustness and fairness in algorithmic decision-making—especially in settings involving skewed or heavy-tailed data.
Appendix A Proofs
Proof of Theorem 1.
We first prove the equivalence statement. Some statements rely on properties of $\mathrm{GD}_{n}$ in Proposition 1 in Section 4, which will be proved later separately.
Sufficiency. By the definition of $i$-th order Gini deviation in (4), we have
$\rho(X)=\sum_{i=1}^{n}a_{i}{\rm GD}_{i}(X)=\mathbb{E}\left[\sum_{i=2}^{n}\frac{a_{i}}{i}\left(\max\{X_{1},\dots,X_{i}\}-\min\{X_{1},\dots,X_{i}\}\right)\right]+\mathbb{E}\left[\frac{a_{1}}{2}|X_{1}-X_{2}|\right].$
This implies that $\rho$ satisfies [A1]. It follows from Proposition 1 that $\rho$ defined by (3) is a signed Choquet integral, and thus, it follows from Theorem 1 of Wang et al. (2020a) that $\rho$ satisfies symmetry, comonotonic additivity, and uniform norm continuity. This completes the proof of sufficiency.
Necessity. According to Theorem 1 in Wang et al. (2020a), any mapping $\rho$ satisfying law invariance (implied by [A1]), [A3] and [A4] can be characterized as a signed Choquet integral with an associated distortion function denoted by $h$. We claim that $h$ is a polynomial of degree at most $n$. To see this, by [A1], there exists a measurable function $f:\mathbb{R}^{n}\to\mathbb{R}$ such that
$\rho(X)=\mathbb{E}[f(X_{1},\dots,X_{n})]=\int_{0}^{\infty}h(\mathbb{P}(X>x))\,\mathrm{d}x+\int_{-\infty}^{0}(h(\mathbb{P}(X>x))-h(1))\,\mathrm{d}x,~~X\in\mathcal{X},$
where $X_{1},\dots,X_{n}$ are iid copies of $X$. Let $X$ be a Bernoulli random variable with $\mathbb{P}(X=0)=1-\mathbb{P}(X=1)=1-t$. Define
$A_{i}=\{\mathbf{z}\in\{0,1\}^{n}:\|\mathbf{z}\|_{1}=i\},~~~~i=0,1,\dots,n,$
where $\|\mathbf{z}\|_{1}$ represents the $1$-norm of $\mathbf{z}$. By straightforward calculation, we have
$h(t)=\rho(X)=\mathbb{E}[f(X_{1},\dots,X_{n})]=\sum_{i=0}^{n}\left(t^{i}(1-t)^{n-i}\sum_{\mathbf{z}\in A_{i}}f(\mathbf{z})\right),$
which means that $h$ is a polynomial of degree at most $n$. Let $(c_{0},c_{1},\dots,c_{n})\in\mathbb{R}^{n+1}$ satisfy
$h(t)=\sum_{i=0}^{n}c_{i}t^{i},~~t\in[0,1].$ (12)
The symmetry [A2] implies that (see Wang et al. (2020a, Proposition 2))
$h(t)=h(1-t)-h(1),~~t\in[0,1].$
This yields $h(0)=h(1)=0$, and thus $c_{0}=0$. Moreover,
$h(t)$ $=\frac{h(t)+h(1-t)}{2}$
$=\frac{1}{2}\sum_{i=1}^{n}c_{i}(t^{i}+(1-t)^{i})=\sum_{i=1}^{n}\left(\frac{-ic_{i}}{2}\right)\left(\frac{1}{i}\right)(1-t^{i}-(1-t)^{i}).$
By the representation result of ${\rm{GD}}_{i}$ in terms of signed Choquet integral in Proposition 1 of Section 4.2, $\rho$ admits the form of (3) with $a_{1}+a_{2}=-c_{2}$ and $a_{i}=-ic_{i}/2$ for $i=3,\dots,n$.
Sufficiency for other axioms. Nonnegativity [A5] follows from (4) and $a_{1},\dots,a_{n}\geq 0$. [A6]–[A7] follow directly from properties of signed Choquet integrals in Wang et al. (2020a). [A8]–[A11] follow the fact that each $\mathrm{GD}_{i}$ has a concave distortion function in Proposition 1 and $a_{1},\dots,a_{n}\geq 0$; see Theorem 3 of Wang et al. (2020a). To show [A12], note that $\mathbb{E}[\max\{X_{1},\dots,X_{n}\}]<\mathbb{E}[X_{1}+\dots+X_{n}]=n\mathbb{E}[X_{1}]$, and hence $\mathrm{GD}_{i}(X)/\mathbb{E}[X]\leq 1$ for all $X\in\mathcal{X}_{+}$. Example 1 below illustrates that the functionals $\mathrm{GD}_{i}(X)/\mathbb{E}[X]$ share the same sequence $(X_{\epsilon})$ of random variables such that $\mathrm{GD}_{i}(X_{\epsilon})/\mathbb{E}[X_{\epsilon}]\to 1$ as $\epsilon\downarrow 0$. Therefore, as their convex combination, $\rho$ satisfies $\rho(X_{\epsilon})/\mathbb{E}[X_{\epsilon}]\to 1$, and hence [A12] holds. ∎
Proof of Theorem 2..
The properties follow from the corresponding properties of $\mathrm{GD}_{n}$ in Theorem 1 and the fact that the ratio of a convex function to a positive affine function is quasi-convex (see e.g., Boyd and Vandenberghe (2004, Section 3.4.4)). ∎
Proof of Proposition 1..
(i) Since
$\mathbb{P}\left(\max\left\{X_{1},\ldots,X_{n}\right\}\leq x\right)=(\mathbb{P}(X\leq x))^{n}=(F_{X}(x))^{n},$
and
$\mathbb{P}\left(\min\left\{X_{1},\ldots,X_{n}\right\}\leq x\right)=1-(\mathbb{P}(X>x))^{n}=1-(1-F_{X}(x))^{n},$
we have
$\mathrm{GD}_{n}(X)$ $=\frac{1}{n}\int_{\mathbb{R}}x\,\mathrm{d}(F_{X}(x)^{n}-1+(1-F_{X}(x))^{n})$ (13)
$=\int_{0}^{1}F^{-1}_{X}(t)(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t.$
(ii) It follows from (13) that $\mathrm{GD}_{n}$ is a signed Choquet integral with a concave distortion function in (6). ∎
Proof of Proposition 2..
For $\mu\in\mathbb{R}$ and $\sigma>0$, consider the set
$\mathcal{M}(\mu,\sigma)=\{X\in L^{2}:\mathbb{E}[X]=\mu,~\mathrm{SD}(X)=\sigma\}.$
Using the Hölder inequality, we obtain, with $h_{n}$ defined in (6),
$\sup_{X\in\mathcal{M}(\mu,\sigma)}\mathrm{GD}_{n}(X)$ $=\sup_{X\in\mathcal{M}(0,\sigma)}\mathrm{GD}_{n}(X)$
$=\sup_{X\in\mathcal{M}(0,\sigma)}\int_{0}^{1}h_{n}^{\prime}(t)F^{-1}_{X}(1-t)\mathrm{d}t$
$\leqslant\sup_{X\in\mathcal{M}(0,\sigma)}\left\|h_{n}^{\prime}\right\|_{2}\left(\int_{0}^{1}\left|F^{-1}_{X}(1-t)\right|^{2}\mathrm{~d}t\right)^{1/2}=\sigma\left\|h_{n}^{\prime}\right\|_{2}.$
Equality is attained if and only if $|F^{-1}_{X}(1-t)|^{2}$ is a multiple of $(h^{\prime}_{n})^{1/2}$, leading to $F^{-1}_{X}(1-t)=h^{\prime}(t)/\left\|h_{n}^{\prime}\right\|_{2}$. Finally, computing the norm gives
$\left\|h_{n}^{\prime}\right\|_{2}=\sqrt{\frac{2}{2n-1}-\frac{2((n-1)!)^{2}}{(2n-1)!}},$
which shows that the upper bound holds and it is sharp. To show that the lower bound is sharp, consider $X_{\epsilon}$ in Example 1, for which we have $\mathrm{GD}_{n}(X_{\epsilon})\leq\epsilon$, and $\mathrm{SD}_{n}(X_{\epsilon})\approx\epsilon^{1/2}$ for $\epsilon>0$ small. ∎
Proof of Proposition 3..
By Proposition 6 in Appendix C, since $n\leq m$, we know $\mathrm{GD}_{n}(X)/\mathrm{GD}_{m}(X)\leq 1$ for $X\in L^{1}_{*}$. For the lower bound, (8) gives, with $h_{n}$ defined in (6),
$\frac{\mathrm{GD}_{n}(X)}{\mathrm{GD}_{m}(X)}\geq\inf_{t\in(0,1)}\frac{h_{n}(t)}{h_{m}(t)}$
By direct computation, we find that this infimum is attained at $t=0.5$, yielding
$\inf_{t\in(0,1)}\frac{h_{n}(t)}{h_{m}(t)}=\frac{h_{n}(0.5)}{h_{m}(0.5)}=\frac{m}{n}\frac{1-2^{1-n}}{1-2^{1-m}}.$
To show that the bound is sharp, consider the two-point distribution
$F=(1-p)\delta_{x}+p\delta_{y},~~~\text{where}~x<y~\text{and}~p\in(0,1),$
where $\delta_{x}$ is the point-mass at $x$. We can compute
$\frac{\mathrm{GD}_{n}(X)}{\mathrm{GD}_{m}(X)}=\frac{h_{n}(p)(y-x)}{h_{m}(p)(y-x)}=\frac{m(1-p^{n}-(1-p)^{n})}{n(1-p^{m}-(1-p)^{m})}.$
Taking the limit as $p\downarrow 0$, we have
$\lim_{p\downarrow 0}\frac{m(1-p^{n}-(1-p)^{n})}{n(1-p^{m}-(1-p)^{m})}=\lim_{p\downarrow 0}\frac{-p^{n-1}+(1-p)^{n-1}}{-p^{m-1}+(1-p)^{m-1}}=1.$
Hence, the upper bound in (9) holds. The lower bound is attained by setting $p=0.5$. ∎
Proof of Theorem 3..
It is straightforward to check
$\operatorname*{arg\,min}_{x\in\mathbb{R}}\int_{\mathbb{R}^{n}}S\left(x,y_{1},\dots,y_{n}\right)\mathrm{d}F(y_{1})\cdots\mathrm{d}F(y_{n})$
$=\operatorname*{arg\,min}_{x\in\mathbb{R}}\mathbb{E}\left[\left(nx-\left(\max\{X_{1},\dots,X_{n}\}-\min\{X_{1},\dots,X_{n}\}\right)\right)^{2}\right]$
$=\frac{1}{n}\mathbb{E}\left[\max\{X_{1},\dots,X_{n}\}-\min\{X_{1},\dots,X_{n}\}\right].$
Therefore, $\mathrm{GD}_{n}$ is $n$-observation $\mathcal{M}^{2}$-elicitable. Using the fact that $S$ and $S^{*}$ differs by a constant term in $x$, we can see that $S^{*}$ also elicits $\mathrm{GD}_{n}$, and it is finite on $\mathcal{M}^{1}$. ∎
Proof of Theorem 4..
Any polynomial $h$ of degree $n$ or less can be written as
$h(t)=a_{n}(1-t)^{n}+a_{n-1}(1-t)^{n-1}+\dots+a_{1}(1-t)+a_{0},~~~t\in[0,1],$
for some $a_{0},a_{1},\dots,a_{n}\in\mathbb{R}$ with $a_{0}=-(a_{1}+a_{2}+\dots+a_{n})$. For iid random variables $X,X_{1},\dots,X_{n}$ with distribtion $F\in\mathcal{M}^{2}$, we have
$\rho(X)=$ $\int_{0}^{1}F_{X}^{-1}(1-t)\,\mathrm{d}h(t)$
$=$ $\int_{0}^{1}F_{X}^{-1}(1-t)\,\mathrm{d}(a_{n}(1-t)^{n}+\dots+a_{1}(1-t)+a_{0})=-\sum_{i=1}^{n}a_{i}\mathbb{E}[\max\{X_{1},\dots,X_{i}\}].$
For the function $S$ defined by
$S\left(x,y_{1},\dots,y_{n}\right)=\left(x+\sum_{i=1}^{n}a_{i}\max\{y_{1},\dots,y_{i}\}\right)^{2},$
it is easy to check
$\operatorname*{arg\,min}_{{x\in\mathbb{R}}}\int_{\mathbb{R}^{n}}S\left(x,\mathbf{y}\right)\mathrm{d}F^{n}(\mathbf{y})=\operatorname*{arg\,min}_{{x\in\mathbb{R}}}\mathbb{E}\left[\left(x+\sum_{i=1}^{n}a_{i}\max\{X_{1},\dots,X_{i}\}\right)^{2}\right]=\rho(X),$
and hence $\rho$ is $n$-observation $\mathcal{M}^{2}$-elicitable. Removing the square terms from $S$ which does not involve $x$, we get a function $S^{*}$ that elicits $\rho$ on $\mathcal{M}^{1}$. ∎
Proof of Proposition 5..
It is straightforward to check
$\operatorname*{arg\,min}_{x\in\mathbb{R}}\int_{\mathbb{R}^{n}}S\left(x,y_{1},\dots,y_{n}\right)\mathrm{d}F(y_{1})\cdots\mathrm{d}F(y_{n})$
$=\operatorname*{arg\,min}_{x\in\mathbb{R}}\mathbb{E}\left[x^{2}X_{1}-\frac{2x}{n}\max\{X_{1},\dots,X_{n}\}-\min\{X_{1},\dots,X_{n}\}\right]$
$=\frac{1}{n}\mathbb{E}\left[\max\{X_{1},\dots,X_{n}\}-\min\{X_{1},\dots,X_{n}\}\right]/\mathbb{E}[X_{1}]=\mathrm{GC}_{n}(X).$
Therefore, $\mathrm{GC}_{n}$ is $n$-observation $\mathcal{M}^{1}$-elicitable. ∎
Proof of Theorem 5.
The Law of Large Numbers yields $\widehat{x}_{N}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\mathbb{E}[X]$. By Theorem 2.6 of Krätschmer et al. (2014), the empirical estimator for a finite convex risk measure on $L^{1}$ is consistent, that is, $\widehat{\mathrm{GD}}_{n}(N)+\widehat{x}_{N}\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\mathrm{GD}_{n}(X)+\mathbb{E}[X]$, and this gives $\widehat{\mathrm{GD}}_{n}(N)\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\mathrm{GD}_{n}(X)$. Moreover, by the Continuous Mapping Theorem, since the mapping $(a,b)\to\frac{a}{b}$ is continuous at $(\mathrm{GD}_{n}(X),\mathbb{E}[X])$ with $\mathbb{E}[X]>0$, we have $\widehat{\mathrm{GC}}_{n}(N)\stackrel{{\scriptstyle\mathbb{P}}}{{\rightarrow}}\mathrm{GC}_{n}(X)$.
Next, we will show the asymptotic normality. Let $B=(B_{t})_{t\in[0,1]}$ be a standard Brownian bridge, and let $d_{N}=\sqrt{N}(\widehat{\mathrm{GD}}_{n}(N)-\mathrm{GD}_{n}(X))$, $e_{N}=\sqrt{N}(\widehat{x}_{N}-\mathbb{E}[X])$, and $g_{N}=\sqrt{N}(\widehat{\mathrm{GC}}_{n}(N)-\mathrm{GC}_{n}(X)))$. Note that $\mathrm{GD}_{n}$ can be written as an integral of the quantile, that is,
$\mathrm{GD}_{n}(X)=\int_{0}^{1}F^{-1}(t)(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t.$
Denote by $A_{N}$ the empirical quantile process, that is,
$A_{N}(t)=\sqrt{N}(\widehat{F}^{-1}_{N}(t)-F^{-1}(t)),~~~t\in(0,1).$
It follows that
$d_{N}=\int_{0}^{1}A_{N}(t)(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t.$
It is well known that, under Assumption 1, as $N\to\infty$, $A_{N}$ converges to the Gaussian process $B/\tilde{f}$ in $L^{\infty}[\delta,1-\delta]$ for any $\delta>0$ (see e.g., Del Barrio et al. (2005)). This yields
$\int_{\delta}^{1-\delta}A_{N}(t)(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t\stackrel{{\scriptstyle\mathrm{d}}}{{\rightarrow}}\int_{\delta}^{1-\delta}\frac{B_{t}}{\tilde{f}(t)}(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t.$
Next, we verify
$\int_{\delta}^{1-\delta}(t^{n-1}-(1-t)^{n-1})\frac{B_{t}}{\tilde{f}(t)}\,\mathrm{d}t\to\int_{0}^{1}\frac{B_{t}}{\tilde{f}(t)}(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t\mbox{~~~as $\delta\downarrow 0$}.$ (14)
Denote by $w_{t}=t(1-t)$. Since $X\in L^{\gamma}$, we have, for some $C>0$,
$|F^{-1}(t)|\leq Cw_{t}^{-1/\gamma}~~\text{and}~\frac{1}{\tilde{f}(t)}=\frac{\,\mathrm{d}F^{-1}(t)}{\,\mathrm{d}t}\leq Cw_{t}^{-1/\gamma-1}.$
Note that $B_{t}=o_{\mathbb{P}}(w_{t}^{1/2-\epsilon})$ for any $\epsilon>0$ as $t\to 0$ or $1$. Hence, for some $\eta>0$,
$\left|\frac{B_{t}}{\tilde{f}(t)}(t^{n-1}-(1-t)^{n-1})\right|=o_{\mathbb{P}}(w_{t}^{\eta-1})\mbox{~~~for $t\in(0,1)$},$
and this guarantees (14). Using the convariance property of the Brownian bridge, that is, $\mathrm{Cov}(B_{t},B_{s})=s-st$ for $s<t$, we have
$\mathrm{Var}\left[\int_{0}^{1}\frac{B_{s}(s^{n-1}-(1-s)^{n-1})}{\tilde{f}(s)}\mathrm{d}s\right]$
$=\mathbb{E}\left[\int_{0}^{1}\int_{0}^{1}\frac{(s^{n-1}-(1-s)^{n-1})(t^{n-1}-(1-t)^{n-1})B_{s}B_{t}}{\tilde{f}(s)\tilde{f}(t)}\mathrm{d}t\mathrm{d}s\right]$
$=\int_{0}^{1}\int_{0}^{1}\frac{(s^{n-1}-(1-s)^{n-1})(t^{n-1}-(1-t)^{n-1})(s\wedge t-st)}{\tilde{f}(s)\tilde{f}(t)}\mathrm{d}t\mathrm{d}s.$
Thus, $\sqrt{N}(\widehat{\mathrm{GD}}_{n}(N)-\mathrm{GD}_{n}(X))\stackrel{{\scriptstyle\mathrm{d}}}{{\rightarrow}}\mathrm{N}(0,\sigma^{2}_{\mathrm{GD}_{n}})$ in which $\sigma^{2}_{\mathrm{GD}_{n}}$ is given by (10).
Note that
$g_{N}=\sqrt{N}\left(\frac{\widehat{\mathrm{GD}}_{n}(N)}{\widehat{x}_{N}}-\frac{\mathrm{GD}_{n}(X)}{\mathbb{E}[X]}\right).$
Since
$d_{N}=\int_{0}^{1}A_{N}(t)(t^{n-1}-(1-t)^{n-1})\,\mathrm{d}t,~~~e_{N}=\int_{0}^{1}A_{N}(t)\,\mathrm{d}t,$
by Delta method, we have
$g_{N}=\textbf{R}^{\top}\textbf{L}+o_{p}(1),$
where
$\textbf{R}=\left(\frac{1}{\mathbb{E}[X]},-\frac{\mathrm{GD}_{n}(X)}{(\mathbb{E}[X])^{2}}\right),~~~~\textbf{L}=(d_{N},e_{N}).$
Since
$(d_{N},e_{N})\stackrel{{\scriptstyle\mathrm{d}}}{{\rightarrow}}\left(\int_{0}^{1}\frac{B_{s}}{\tilde{f}(s)}(s^{n-1}-(1-s)^{n-1})\mathrm{d}s,\int_{0}^{1}\frac{B_{s}}{\tilde{f}(s)}\mathrm{d}s\right),$
we have
$g_{N}\stackrel{{\scriptstyle\mathrm{d}}}{{\rightarrow}}\frac{1}{\mathbb{E}[X]}\int_{0}^{1}\frac{B_{s}}{\tilde{f}(s)}\left(s^{n-1}-(1-s)^{n-1}-\mathrm{GC}_{n}(X)\right)\mathrm{d}s.$
Using the convariance property of the Brownian bridge, we have
$\mathrm{Var}\left[\frac{1}{\mathbb{E}[X]}\int_{0}^{1}\frac{B_{s}}{\tilde{f}(s)}\left(s^{n-1}-(1-s)^{n-1}-\mathrm{GC}_{n}(X)\right)\mathrm{d}s\right]$
$=\mathbb{E}\left[\int_{0}^{1}\int_{0}^{1}\frac{(s^{n-1}-(1-s)^{n-1}-\mathrm{GC}_{n}(X))(t^{n-1}-(1-t)^{n-1}-\mathrm{GC}_{n}(X))B_{s}B_{t}}{(\mathbb{E}[X])^{2}\tilde{f}(s)\tilde{f}(t)}\mathrm{d}t\mathrm{d}s\right]$
$=\int_{0}^{1}\int_{0}^{1}\frac{(s^{n-1}-(1-s)^{n-1}-\mathrm{GC}_{n}(X))(t^{n-1}-(1-t)^{n-1}-\mathrm{GC}_{n}(X))(s\wedge t-st)}{(\mathbb{E}[X])^{2}\tilde{f}(s)\tilde{f}(t)}\mathrm{d}t\mathrm{d}s.$
Thus, $\sqrt{N}(\widehat{\mathrm{GC}}_{n}(N)-\mathrm{GC}_{n}(X))\stackrel{{\scriptstyle\mathrm{d}}}{{\rightarrow}}\mathrm{N}(0,\sigma^{2}_{\mathrm{GC}_{n}})$ in which $\sigma^{2}_{\mathrm{GC}_{n}}$ is given by (11). ∎
Appendix B Explicit GD and GC formulas for parametric distributions
In this section, we derive explicit formulas or integral representations for $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ for several commonly used parametric distributions, including Bernoulli, Beta, log-normal, exponential, and Pareto distributions.
  • (i)
    Bernoulli distribution: Let $X\sim\mathrm{B}(p)$, where $p\in(0,1)$. Then
    $\mathrm{GD}_{n}(X)=\int_{1-p}^{1}(t^{n-1}-(1-t)^{n-1})\mathrm{d}t=\frac{1}{n}(1-p^{n}-(1-p)^{n}).$
    Since $\mathbb{E}[X]=p$, giving
    $\mathrm{GC}_{n}(X)=\frac{1-p^{n}-(1-p)^{n}}{np}.$
    Figure 13 presents the curves of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ as $p$ and $n$ vary.
    Figure 13: Values of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ for $X\sim\mathrm{B}(p)$ with $p\in(0,1)$
  • (ii)
    Beta distribution: For a Beta-distributed random variable $X\sim\text{Beta}(a,b)$ with shape parameters $a,b>0$, the CDF is given by the regularized incomplete Beta function:
    $F_{X}(x)=I_{x}(a,b)=\frac{\int_{0}^{x}u^{a-1}(1-u)^{b-1}\mathrm{d}u}{B(a,b)},$
    where $B(a,b)=\int_{0}^{1}u^{a-1}(1-u)^{b-1}\mathrm{d}u$ is the Beta function. The quantile function is denoted as: $F_{X}^{-1}(t)=I_{t}^{-1}(a,b),$ where $I_{t}^{-1}(a,b)$ is the inverse regularized incomplete Beta function. Thus, $\mathrm{GD}_{n}(X)$ for the Beta distribution is:
    $\mathrm{GD}_{n}(X)=\int_{0}^{1}I_{t}^{-1}(a,b)\left(t^{n-1}-(1-t)^{n-1}\right)\mathrm{d}t.$
    Since $\mathbb{E}[X]=a/(a+b)$, we obtain
    $\mathrm{GC}_{n}(X)=\frac{\mathrm{GD}_{n}(X)(a+b)}{a}.$
    Figures 14 and 15 present the curves of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ as $a$ $b$, and $n$ vary.
    Figure 14: Values of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ for $X\sim\text{Beta}(a,1)$ with $a\in(0,5)$
    Figure 15: Values of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ for $X\sim\text{Beta}(1,b)$ with $b\in(0,10)$
  • (iii)
    Log-Normal distribution: For a log-normal random variable $X\sim\mathrm{LN}(\mu,\sigma^{2})$, the CDF is given by
    $F_{X}(x)=\Phi\left(\frac{\ln x-\mu}{\sigma}\right),$
    where $\Phi$ is the standard normal CDF. The quantile function is given by
    $F_{X}^{-1}(t)=\exp(\mu+\sigma\Phi^{-1}(t)).$
    Thus, $\mathrm{GD}_{n}(X)$ is
    $\mathrm{GD}_{n}(X)=\int_{0}^{1}\exp(\mu+\sigma\Phi^{-1}(t))\left(t^{n-1}-(1-t)^{n-1}\right)\mathrm{d}t.$
    The expectation of $X$ is $\mathbb{E}[X]=e^{\mu+\sigma^{2}/2}$, giving
    $\mathrm{GC}_{n}(X)=\frac{\mathrm{GD}_{n}(X)}{e^{\mu+\sigma^{2}/2}}=\int_{0}^{1}\exp\left(\sigma\Phi^{-1}(t)-\frac{\sigma^{2}}{2}\right)\left(t^{n-1}-(1-t)^{n-1}\right)\mathrm{d}t,$
    which is independent in $\mu$. Figures 16 and 17 present the curves of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ as $\mu$ $\sigma$, and $n$ vary.
    Figure 16: Values of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ for $X\sim\mathrm{LN}(\mu,1)$ with $\mu\in(-3,3)$
    Figure 17: Values of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ for $X\sim\mathrm{LN}(1,\sigma^{2})$ with $\sigma\in(0,2)$
  • (iv)
    Exponential distribution: For an exponential random variable $X\sim\text{exp}(\lambda)$, the CDF is given by $F_{X}(x)=1-e^{-\lambda x},~x\geq 0.$ The quantile function is
    $F_{X}^{-1}(t)=-\frac{1}{\lambda}\ln(1-t),\quad t\in[0,1).$
    Thus, we have
    $\mathrm{GD}_{n}(X)$ $=\int_{0}^{1}\left(-\frac{1}{\lambda}\ln(1-t)\right)\left(t^{n-1}-(1-t)^{n-1}\right)\,\mathrm{d}t$
    $=\frac{1}{n\lambda}\sum_{k=1}^{n-1}\frac{1}{k}.$
    The expectation of $X$ is $\mathbb{E}[X]=1/\lambda$, giving
    $\mathrm{GC}_{n}(X)=\frac{1}{n}\sum_{k=1}^{n-1}\frac{1}{k},$
    which is independent in $\lambda$. Figure 18 presents the curves of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ as $\lambda$ and $n$ vary.
    Figure 18: Values of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ for $X\sim\text{exp}(\lambda)$ with $\lambda\in(0,2)$
  • (v)
    Pareto distribution: For a Pareto-distributed random variable $X\sim\text{Pareto}(\alpha,x_{m})$, the CDF is given by
    $F_{X}(x)=1-\left(\frac{x_{m}}{x}\right)^{\alpha},\quad x\geq x_{m}.$
    The quantile function is
    $F_{X}^{-1}(t)=x_{m}(1-t)^{-1/\alpha},\quad t\in[0,1).$
    Thus, we have
    $\mathrm{GD}_{n}(X)$ $=\int_{0}^{1}\left(x_{m}(1-t)^{-1/\alpha}\right)\left(t^{n-1}-(1-t)^{n-1}\right)\,\mathrm{d}t$
    $=x_{m}\left(B(n,1-1/\alpha)-\frac{1}{n-1/\alpha}\right).$
    The expectation of $X$ is $\mathbb{E}[X]=\alpha x_{m}/(\alpha-1)$ for $\alpha>1$ (for $\alpha\leq 1$, $\mathbb{E}[Y]$ is infinite), giving
    $\mathrm{GC}_{n}(X)=\frac{\alpha-1}{\alpha}\left(B(n,1-1/\alpha)-\frac{1}{n-1/\alpha}\right),$
    where $B(n,1-1/\alpha)$ is the Beta function. It is clear that $\mathrm{GC}_{n}(X)$ is independent in $x_{m}$. Figures 19 and 20 present the curves of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ as $\alpha$ and $n$ vary.
    Figure 19: Values of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ for $X\sim\text{Pareto}(\alpha,1)$ with $\alpha\in(1,5)$
    Figure 20: Values of $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ for $X\sim\text{Pareto}(2,x_{m})$ with $x_{m}\in(0,5)$
Appendix C Monotonicity with respect to the order
In this section, we discuss how $\mathrm{GD}_{n}$ and $\mathrm{GC}_{n}$ changes as the parameter $n$ varies. Recall that $\mathrm{GD}_{3}=\mathrm{GD}_{2}$. The situation for $n\geq 4$ is however different. We get from (6),
$h_{4}(t)=\frac{1}{2}t(1-t)(2-t+t^{2})\leq h_{3}(t)=h_{2}(t).$
Therefore, $\mathrm{GD}_{4}\leq\mathrm{GD}_{3}=\mathrm{GD}_{2}$, and $\mathrm{GD}_{4}\neq\mathrm{GD}_{3}$. On the other hand, since $2-t+t^{2}\geq 7/4$ for $t\in[0,1]$, we have $\mathrm{GD}_{4}(X)/\mathrm{GD}_{3}(X)\in[7/8,1]$ whenever $\mathrm{GD}_{3}(X)\neq 0$ (see also Proposition 3 below). In the following proposition, we can show that $\mathrm{GD}_{n}$ decreases in $n$, and this result was obtained by Gavilan-Ruiz et al. (2024) with a different proof.
Proposition 6.
For $X\in L^{1}$ (resp. $X\in L^{1}_{+}$), $\mathrm{GD}_{n}(X)$ (resp. $\mathrm{GC}_{n}(X)$) decreases to 0 as $n\to\infty$.
Proof.
We only prove the statement for $\mathrm{GD}_{n}$, as the proof for $\mathrm{GC}_{n}$ is identical. Let $n\geq 2$ and $X_{1},\dots,X_{n+1}$ be iid copies of $X$. Define $A_{k}=\max\left\{X_{1},\dots X_{k}\right\}$ and $B_{k}=\min\left\{X_{1},\dots X_{k}\right\}$ for $k\in[n+1]$. By symmetry, for each $i\in[n+1]$, we have
$\mathbb{E}\left[A_{n}-B_{n}\mid A_{n+1},B_{n+1}\right]=\mathbb{E}\left[\max_{j\in[n+1]\setminus\{i\}}X_{j}-\min_{j\in[n+1]\setminus\{i\}}X_{j}\mid A_{n+1},B_{n+1}\right].$
By setting $R_{n}=A_{n}-B_{n}$, we have
$\sum_{i=1}^{n+1}\mathbb{E}\left[\max_{j\in[n+1]\setminus\{i\}}X_{j}-\min_{j\in[n+1]\setminus\{i\}}X_{j}\mid A_{n+1},B_{n+1}\right]=\mathbb{E}[(n+1)R_{n}\mid A_{n+1},B_{n+1}].$
Note that
$\sum_{i=1}^{n+1}\max_{j\in[n+1]\setminus\{i\}}X_{j}=nA_{n+1}+X_{n:n+1}\mbox{~~~and~~~}\sum_{i=1}^{n+1}\max_{j\in[n+1]\setminus\{i\}}X_{j}=nB_{n+1}+X_{2:n+1},$
where $X_{n:n+1}$ (resp. $X_{2:n+1}$) is the second largest (resp. second smallest) value in $X_{1},\dots,X_{n+1}$. Since $X_{n:n+1}\geq X_{2:n+1}$, We have $\mathbb{E}\left[(n+1)R_{n}\mid A_{n+1},B_{n+1}\right]\geq\mathbb{E}\left[nR_{n+1}\mid A_{n+1},B_{n+1}\right]=nR_{n+1}$. It then follows that
$\mathbb{E}\left[\left.\frac{1}{n}R_{n}\,\right\rvert\,A_{n+1},B_{n+1}\right]\geq\frac{1}{n+1}R_{n+1},$
and thus
$\frac{1}{n}\mathbb{E}[R_{n}]\geq\frac{1}{n+1}\mathbb{E}[R_{n+1}],$
which completes the proof of the first statement. Furthermore, by the dominated convergence theorem, we have
$\lim_{n\to\infty}\mathrm{GD}_{n}(X)=\int_{0}^{1}F_{X}^{-1}(t)\lim_{n\to\infty}\left(t^{n-1}-(1-t)^{n-1}\right)\,\mathrm{d}t=0,$
showing the second statement. ∎
In the following example, we compare the behavior of $\mathrm{GC}$ and $\mathrm{GC}_{n}$ for two groups of different distributions: (i) $X\sim\mathrm{LN}(\mu,\sigma)$ and $Y\sim\text{Pareto}(\alpha,x_{m})$, and (ii) $X\sim\text{Beta}(a,b)$ and $Y\sim\text{Pareto}(\alpha,x_{m})$. The specific distribution details and the expression for computing $\mathrm{GC}_{n}$ are provided in Appendix B. Note that $\mathrm{GC}_{n}$ is independent of $\mu$ for the log-normal distribution and independent of $x_{m}$ for the Pareto distribution.
Figure 21: Survival functions (left panel) and values of $\mathrm{GC}_{n}$ (right panel) for $\mathrm{LN}(0,\sigma)$ and $\mathrm{Pareto}(\alpha,1)$
Figure 22: Survival functions (left panel) and values of $\mathrm{GC}_{n}$ (right panel) for $\mathrm{Beta}(a,2)$ and $\mathrm{Pareto}(\alpha,0.1)$
On the left panel of Figure 21, we show the survival functions of the Pareto and log-normal distributions. Clearly, the Pareto distribution exhibits a slower tail decay, meaning it has a heavier tail compared to the log-normal distribution. On the right panel of Figure 21, we illustrate $\mathrm{GC}_{n}$ as a function of $n$ for both distributions. Our findings reveal an interesting pattern: increasing $n$ enhances the sensitivity of $\mathrm{GC}_{n}$ to tail risk, capturing the disparity between distributions with different tail behaviors. Specifically, when $n=2$, $\mathrm{GC}_{n}$ for the Pareto distribution is lower than that of the log-normal distribution. However, as $n$ increases, $\mathrm{GC}_{n}$ for the Pareto distribution eventually surpasses that of the log-normal distribution. This behavior can be understood from the definition of $\mathrm{GC}_{n}$. The term $\mathrm{GD}_{n}(X)$ is influenced by higher-order differences in the distribution function, making it more sensitive to tail behavior as $n$ increases. Since the Pareto distribution has a heavier tail than the log-normal distribution, the contribution of extreme values becomes more pronounced for larger $n$, leading to a higher $\mathrm{GC}_{n}$. The same phenomenon also occurs for the Pareto and Beta distributions in Figure 22.
From these numerical results, we can see that for a moderate choice of $n$ such as $n\leq 20$, the value of $\mathrm{GC}_{n}(X)$ is typically not close to 0, so that the limiting behavior $\mathrm{GC}_{n}(X)\to 0$ in Proposition 6 does not hurt the practical applicability of $\mathrm{GC}_{n}$. Also note that by combining Theorem 1 and Proposition 6, we have two asymptotic results:
$\lim_{n\to\infty}\sup_{X\in L^{1}_{+}}\mathrm{GC}_{n}(X)=1~~\mbox{ and }~~\sup_{X\in L^{1}_{+}}\lim_{n\to\infty}\mathrm{GC}_{n}(X)=0.$
That is, the limit and the supremum cannot be interchanged. Even for large values of $n$, there exists some random variables in $L^{1}_{+}$ for which $\mathrm{GC}_{n}$ is not close to 0, and the index $\mathrm{GC}_{n}$ can be used to analyze the inequality in some extreme distributions.